Probability of a number greater than 7, P(>7) = 3/10, Probability of an even number, P(even number) = 5/10 = 50%, Probability of an 8., P(8) = 1/10 = 10%, Probability of landing on a 4 or a 5, P(4 or 5) = 2/10=20%, Probability of orange or an odd number, P(orange or odd) = 6/10 = 60%, Probability of picking a green marble, replacing it, and then picking a red marble, 4/10 * 3/10 = 12%, The probability of picking a red marble, keeping it, and then green marble, 3/10 *4/9 = 12/90 = 4/30 = 13%, The probability of picking a red ball, replacing it, then picking a blue ball., 3/10 * 2/10 = 6/100 = 6%, The probability of picking a red ball, keeping it, then picking a blue ball., 3/10 * 2/9 = 2/30 = 1/15 = 6.7%
0%
Probability Flip
共有
共有
Emeliesutter
さんの投稿です
G7
Math
Probability
コンテンツの編集
セーブ
いいね!
印刷
埋め込み
もっと見る
割り当て
リーダーボード
もっと表示する
表示を少なくする
このリーダーボードは現在非公開です。公開するには
共有
をクリックしてください。
このリーダーボードは、リソースの所有者によって無効にされています。
このリーダーボードは、あなたのオプションがリソースオーナーと異なるため、無効になっています。
オプションを元に戻す
タイルをめくる
は自由形式のテンプレートです。リーダーボード用のスコアは生成されません。
ログインが必要です
表示スタイル
フォント
サブスクリプションが必要です
オプション
テンプレートを切り替える
すべてを表示
オープン結果
リンクをコピー
QRコード
削除
自動保存:
を復元しますか?